OpenStudy (anonymous):

differentiate: f(x) = (x^4+5x^4)^-1

OpenStudy (anonymous):

I'll show you how to do it.

OpenStudy (anonymous):

This involves the chain rule.

OpenStudy (anonymous):

So, if x is inside of some other type of function, you must take the derivatives of what happens to x from inside of the function all the way out. In this case, x becomes the expression (x^4 + 5x^4), and then it enters a function which raises that expression to the power of -1. So we must take the derivative of the expression that x becomes, and multiply it by the derivative of the function that it gets put into (the one that gives it a -1 exponent). This gives the following: \[\frac{d}{dx}(x^4+5x^4)^{-1} = \frac{d}{dx}(x^4+5x^4)\frac{d}{du}(u^{-1})\] where u is the expression (x^4+5x^4)

OpenStudy (anonymous):

So this becomes \[(4x^3+20x^3)(-u^{-2})\]

OpenStudy (anonymous):

Since u is equal to \[x^4+5x^4\] the final answer is \[(4x^3+20x^3)(-(x^4+5x^4)^{-2})=-\frac{4x^3+20x^3}{(x^4+5x^4)^{2}}\]

OpenStudy (anonymous):

Which you could simplify further if you wanted

OpenStudy (anonymous):

Thank you very much

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

did you finish simplifying it?

OpenStudy (anonymous):

\[-\frac{24x^3}{36x^8}=-\frac{2}{3x^5}\]